OpenStudy (anonymous):

trig help PLEASE!! If cos A = -4/5 for angle A in Quadrant II, find cos A/2.

OpenStudy (anonymous):

if you look the inverse of cos(-4/5 )you get A= 143.13 deg. A/2=143.13/2= 71.56deg. cos 71.56 =approx.3/10